The mean free path of CO2 molecules in a vessel is 790 Å, density is 1.7 kg/m3.Calculate the diameter of CO2 molecule


1
Expert's answer
2021-10-01T12:18:36-0400

The mean free path can be expressed in terms of pressure and diameter:


"l=\frac{k_BT}{\sqrt2\pi d^2p},\\\space\\
d=\sqrt{\frac{k_BT}{l\pi p\sqrt2}}=3.4\text{Å}."

We took the temperature of 300K and pressure of 101325 Pa.